OpenStudy (pulsified333):

Differentiate the following function: y=7K+Me^t, where K and M are given constants. Enter the derivative here:

OpenStudy (pulsified333):

@greatlife44

OpenStudy (pulsified333):

hi again :D

OpenStudy (greatlife44):

hi :)

OpenStudy (greatlife44):

\[y = 7K + Me^{t}\] so remember, the derivative of a constant is just going to be zero, because think about it there isn't any variable so the slope of a constant is going to be zero \[\frac{ d }{ dx }c = 0 \]

OpenStudy (greatlife44):

now let's go on to this. the derivative of e^{t} is going to just be e^{t} \[\frac{ d }{ dx } e^{t} = e^{t}\] now let's apply these rules to our question. \[\frac{ dy }{ dt }7K + \frac{ dy }{ dt } Me^{t} = 0 + Me^{t}\] \[\frac{ dy }{ dt } = Me^{t}\]

OpenStudy (pulsified333):

but why wouldn't M be zero as well

OpenStudy (greatlife44):

so when we say \[\frac{ dy }{ dt }\] this means we are taking the derivative of y with respect to t

OpenStudy (pulsified333):

oh okay

OpenStudy (greatlife44):

think of it like this \[\frac{ dy }{ dt } Me^{t} = M \frac{ dy }{ dt }e^{t}\] remember the derivative of e^{t} is just going to be e^{t}

OpenStudy (greatlife44):

@Pulsified333 try this find the derivative of \[2x \]

OpenStudy (pulsified333):

2

OpenStudy (greatlife44):

okay, 

OpenStudy (greatlife44):

so we know that the slope of a line is going to be constant

OpenStudy (pulsified333):

right

OpenStudy (greatlife44):

 y = 2 what about the derivative of this?

OpenStudy (greatlife44):

that's not the same as if we took the derivative of say \[y = 2e^{x}\] in this case 2 is your constant.

OpenStudy (greatlife44):

because remember if 2 which is our constant = 0 then the derivative would be zero because remember \[\frac{ d }{ dx } 2e^{x} = 2e^{x}, 2*e^{x}, so~our~constant~can't~be~ zero. \]

OpenStudy (pulsified333):

true

OpenStudy (greatlife44):

think about it @Pulsified333 \[\frac{ d }{ dx } Ae^{x} = Ae^{x}\]

OpenStudy (greatlife44):

in this case our constant is not zero, because then it would mean 0*e^{x} which would just be zero. I guess there's a better way to explain it but yeah I guess that's the idea.

OpenStudy (pulsified333):

I get it now thank you :D

OpenStudy (greatlife44):

great :)